# Short vs. Long-term Coordination of Drones: When Distributed Optimization Meets Deep Reinforcement Learning

### v2

@@ III-B Problem Formulation
$\mathsf{Eff}(\mathcal{A}^{u},t)=\frac{1}{U}\cdot\frac{\sum^{N}_{n=1}\sum^{S}_{s=1}V^{\prime}_{n,s}(t)}{\sum^{N}_{n=1}\sum^{S}_{s=1}V_{n,s}(t)}.$ (1)
@@
$E(\mathcal{A}^{u},t)=P^{\mathsf{f}}_{u}\cdot t^{\mathsf{f}}(\mathcal{A}^{u})+P^{\mathsf{h}}_{u}\cdot t^{\mathsf{h}}(\mathcal{A}^{u}),$ (3)
where $P^{\mathsf{f}}_{u}$ and $P^{\mathsf{h}}_{u}$ are the flying and hovering power consumption of drone $u$ respectively; $t^{\mathsf{f}}$ and $t^{\mathsf{h}}$ are the flying and hovering time respectively, which are determined by the sensing plans $\mathcal{A}^{u}$ during the period $t$.
@@
$\mathop{\arg\max}\limits_{\mathcal{A}^{u},u\in\mathcal{U}}\;\sum^{U}_{u=1}\sum^{T}_{t=1}[\mathsf{Eff}(\mathcal{A}^{u},t)\cdot\mathsf{Acc}(\mathcal{A}^{u},t)-E(\mathcal{A}^{u},t)].$ (4)
III-C Framework Overview
For high-quality sensing performance, the proposed approach relies on both decentralized drone coordination for short-term navigation and sensing optimization, as well as long-term scheduling of flight directions to adapt to changing environments. Fig.1 illustrates the designed system framework of DO-RL, consisting of four main components:
@@ III-C Framework Overview
Plan generation. This component generates the navigation and sensing options for drones. Given the sensing map, each drone, controlled by a local software agent, autonomously generates a finite number of discrete plans to initialize the overall process. This provides flexibility for the drones at the next stage to choose in a coordinated way. The generated plans are grouped into different flying directions. This initialization stage is executed only once for each drone.
Short-term plan selection. This distributed component coordinates drones to locally select the optimal navigation and sensing options from their generated plans within a period. It outputs and disseminates the aggregated plans $\mathcal{A}(t)$ to all drones in a peer-to-peer way.
DRL-based long-term scheduling. This component is the core of the framework, which leverages the DRL algorithm to enable drones to take their period-by-period actions $a^{u}$ (N, E, S, W, NE, SE, SW, NW or O) based on the distribution of sensor data values. Specifically, drones determine their flying directions with the highest cumulative overall performance of sensing. Then, the generated subset of plans with the corresponding navigation action are chosen as input for the short-term plan selection. Note that drones return to the nearest charging stations after navigation and sensing, and depart from them in the next period.
Periodic state update. This component updates the state of drones (i.e., the current location $c^{u}(t)$ and energy cost $e^{u}(t)$) and the system (i.e., the target $\mathcal{R}(t)$ and required sensor data values $V(t)$) for the long-term scheduling in the next period.
IV Detailed Framework Design
@@ IV-A Plan Generation
Each drone generates nine groups of plans, each associated with an action for a flying direction. These groups consist of multiple individual plans of the same size. A group of plans generated by drone $u$ during a period is defined as $G^{a^{u}}(t)$, where each plan in the group has the same action. A plan with index $l$ is denoted as $\mathcal{A}^{u}_{l}(t)$, $l\leq L$, $\mathcal{A}^{u}_{l}(t)\in G^{a^{u}}(t)$, where $L$ denotes the total number of plans in a group. Therein, each plan has sensing information about the path specifications of the drone, such as the visited cells, the timing of data acquisition over these cells, and the corresponding (normalized) energy consumption. Thus, we define $\mathcal{A}^{u}_{l}(t)=\{\mathcal{A}^{u}(t)|E(\mathcal{A}^{u})\}$, where $\mathcal{A}^{u}(t)$ indicates the visited cells and hovering time; $E(\mathcal{A}^{u})$ denotes the energy consumed by $u$ traveling over these visited cells, i.e., the cost of the plan.
@@
Path finding and energy calculation. A drone generates a plan by determining its path and calculating the energy required for traversing that path, as shown in plan generation part of Fig. 1. As an example of a plan $\mathcal{A}^{u}_{l}(t)\in G^{a^{u}}(t)$, the drone $u$ firstly finds a range to cells $K^{a^{u}}$, which covers the cells close to its location and within its flying direction. Next, it selects a total number of $J^{u}$ visited cells randomly from this range. Then, $u$ finds the shortest possible path among these $J^{u}$ cells via the Dijkstra’s algorithm for the Traveling Salesmen Problem. The path includes the information of which cell and which timeslot the drone uses to collect data, and thus both the traveling time $t^{\mathsf{f}}(\mathcal{A}^{u})$ and hovering time $t^{\mathsf{h}}(\mathcal{A}^{u})$ are calculated. The total energy consumption $E(\mathcal{A}^{u})$ when $u$ travels over the path is then calculated via Eq.(3).
IV-B Short-term Plan Selection
@@
Given the plans generated within a period, drones can adapt their assignments by selecting appropriate plans in response to changes in task targets, enhancing their flexibility and adaptability in dynamic environments. In plan selection, the agent of each drone connects with other agents into a tree communication structure, as shown in Fig. 1, within which it interacts with its children and parent in a bottom-up and top-down fashion to improve plan selections iteratively [8]. Specifically, each agent obtains the aggregated choices from other agents, and chooses one of the plans $\mathcal{A}^{u}_{l}(t)$ such that all choices together $\mathcal{A}(t)=\sum_{u}\mathcal{A}^{u}(t)$ add up to match a target $\mathcal{R}(t)$. The target is used to steer each drone to sense over a unique cell during a timeslot, thereby avoiding the simultaneous sensing of multiple drones within the same cell, i.e., preventing over-sensing and under-sensing.
@@
The purpose of coordination is to match the aggregated plans $\mathcal{A}(t)=\{p_{n,s}(t)\}$ to the target $\mathcal{R}(t)=\{r_{n,s}(t)\}$. Two opposing objectives are optimized: the average difference between $\mathcal{A}(t)$ and $\mathcal{R}(t)$, as well as the energy cost $E(\mathcal{A}^{u})$ of the plan selected by the drone. The cost function for each agent is formulated using the root mean square error (RMSE):
@@
$\begin{split}\mathop{\min}\limits_{\mathcal{A}^{u},u\in\mathcal{U}}\;&(1-\beta)\cdot\sqrt{\frac{\sum^{N}_{n=1}\sum^{S}_{s=1}(p_{n,s}(t)-r_{n,s}(t))^{2}}{N\cdot S}}\\ &+\beta\cdot E(\mathcal{A}^{u}),\end{split}$ (5)
where $\beta$ represents the behavior of an agent. As the value of $\beta$ increases, the agent becomes increasingly selfish, prioritizing plans with lower energy costs at the expense of higher root mean squared error. Each agent (or drone) aims to minimize the system-wide cost of sensing while considering its energy consumption. However, drones have no knowledge about the expected sensing values, e.g., the increase of traffic flow. In the proposed approach, long-term actions enable drones to prioritize for areas with abundant sensor data, enhancing their efficiency in navigation, sensing and recharging.
IV-C DRL-based Long-term Scheduling
The process of determining actions for drones is hindered by what is commonly known as the "dimensional curse." This dimensionality arises exponentially with the increase in the number of drones, cells and time periods, which escalates the complexity of the scheduling problem. Furthermore, the complex distribution of sensor data values on the map further enhances the difficulty in scheduling effectively the actions for the drones. To tackle these issues, the proposed method incorporates the deep reinforcement learning algorithm.
DRL Modeling. We model the problem using state, action, and reward concepts: (i) State: The state $\textbf{s}_{t}$ at period $t$ consists of four components $(\mathcal{S}_{1},\mathcal{S}_{2},\mathcal{S}_{3},\mathcal{S}_{4})$, where $\mathcal{S}_{1}=\{c^{1}(t),...,c^{U}(t)\}$ is the current locations of all drones. Since drones charge at charging stations before taking actions, the location can serve as the index of the charging station, $c^{u}(t)=m$. $\mathcal{S}_{2}=\{e^{1}(t),...,e^{U}(t)\}$ is the current energy consumption of all drones. $\mathcal{S}_{3}=\{p_{1,1}(t),...,p_{n,s}(t),...,p_{N,S}(t)\}$ is the aggregated plans of all drones. $\mathcal{S}_{4}=\{V_{1,1}(t),...,V_{n,s}(t),...,V_{N,S}(t)\}$ is the required distribution of sensor data values. (ii) Action: The action $\textbf{a}_{t}=\{a^{1}_{t},...,a^{U}_{t}\}$ at period $t$ consists of flying directions of all drones. (iii) Reward: Based on the objective function of Eq.(4), the expected immediate local reward of one agent at period $t$ is defined as follows: $\textbf{r}^{u}_{t}=\mathsf{Eff}(t)\cdot\mathsf{Acc}(t)-E^{u}(t)$, where the first part includes the mission efficiency and sensing accuracy calculated via Eq.(1) and (2), while the second part is the energy cost of drone $u$ calculated via Eq.(3).
Global observation. To mitigate the expanded state space and achieve more effective collaborative scheduling, we design a global observation where each drone evaluates an immediate reward based on the aggregated information of all drones. Using the tree communication structure built for the distributed optimization, the agents of the drones aggregate and share their states in a bottom-up fashion, including the aggregated plans and the required sensor data values. They obtain the shared states in the top-down process [8]. The advantages of the global observation are listed as follows: (i) Due to the tree structure, the coordination among agents is highly-efficient with low computational and communication cost [8]. Agents can be self-organized within the tree to make the system more resilient to attacks or failures [12, 11] (ii) Since agents share the aggregated states instead of the agents’ one, the system is scalable and designed to preserve privacy by design [^2] Additional privacy protection mechanisms such as differential privacy and homomorphic encryption apply here [35].. (iii) Each agent is flexible and can choose its own path to navigate among sensing cells and charging stations without over-relying on the policies of other agents.
@@ IV-C DRL-based Long-term Scheduling
Centralized critic multiple actors. Our approach provides a generalized model for multi-agent reinforcement learning with one centralized critic and multiple actors. The actors allow drones to work independently and choose actions based on their observations and policies, allowing for parallel data collection and processing. The centralized critic, however, evaluates the joint actions of drones and guide them toward collectively optimal decisions. As a result, this design enhances the model robustness, adaptability, and coordination capabilities in complex environments. In addition, the proposed method employ Proximal Policy Optimization (PPO) to prevent detrimental updates and improving the stability of the learning process [36].
How to operationalize the solution. Assuming a central authority that guides iterative flying directions of drones using deep reinforcement learning, it directs them to return to and depart from specific charging stations for consecutive flights. These drones, however, belong to various entities, such as aviation agencies, private citizens, and companies, collaborating to create shared navigation and sensing strategies. Note that these strategies rely on the collective sum of all plans rather than individual private sensing information. As a result, the governing authority does not exert any influence over the decision-making process regarding the sensing behavior of the drones. This approach exhibits significant benefits, particularly in terms of privacy preservation and resilience, guaranteeing operational integrity in case of an authority attack.
@@
Overall framework. Fig. 2 illustrates the overall framework of the DRL-based long-term scheduling. The process primarily involves the following steps: Firstly, each agent takes an action for a flying direction based on the global observation $\textbf{o}^{u}_{t}$ shared via coordination as well as the reward $\textbf{r}^{u}_{t}$. This observation contains information of the drone: its current location, energy consumption, the aggregated plans and the required sensor data values. Once all actions are determined, drones select their plans and transition to a new state. Subsequently, the buffer, a data storage structure used for experience replay, stores all transitions of each agent $(\textbf{o}^{u}_{t},\textbf{a}^{u}_{t},\textbf{r}^{u}_{t},\textbf{o}^{u}_{t+1})$. Several groups of transitions $(\textbf{o}^{u}_{k},\textbf{a}^{u}_{k},\textbf{r}^{u}_{k},\textbf{o}^{u}_{k+1})$ are sampled randomly ($H$ groups of transitions) for updating the parameters of both the critic and actor networks. The algorithm is typically an extension of the actor-critic policy gradient approach, employing two deep neural networks for each agent: a critic network $Q(\cdot)$ and an actor network $\pi(\cdot)$. The critic network estimates the reward associated with a transition using Bellman equation, and its parameter $\theta^{Q}$ is then updated by minimizing a loss function $L(\theta^{Q})$:
@@
$\hat{A}^{u}_{k}=\textbf{r}_{k}+\gamma\cdot Q(\textbf{o}^{u}_{k+1},a^{u}_{k+1})-Q(\textbf{o}^{u}_{k},a^{u}_{k}),$ (7)
@@
IV-D Periodic State Update
Location and energy. Each drone changes its location at the next period $t+1$ as $c^{u}_{t+1}=a^{u}_{t}$. Given the selected plan $\mathcal{A}^{u}$, $u$ updates its current energy consumption as $e^{u}_{t+1}=E(\mathcal{A}^{u},t)$.
Predicted sensor data. Due to evolving environmental conditions, drones have no knowledge about the required values of sensor data in the next period. Thus, the predicted sensor data $\hat{V}$ is used to estimate the required one $V$ based on historical information. The $\hat{V}(t)$ at period $t$ is updated in a time-reverse decay, formulated as follows:
$\hat{V}_{n,s}(t)=\sum^{t}_{t^{\prime}=1}(T-t+t^{\prime})\cdot\omega_{n,s}(t^{\prime})\cdot V^{\prime}_{n,s}(t^{\prime}):=V_{n,s}(t),$ (9)
where $V^{\prime}_{n,s}$ denotes the data values collected by drones once they execute their selected plans (the total collected values are aggregated and shared to each agent via the top-down interactions in the distributed optimization process [8]); $\omega_{n,s}(t^{\prime})$ is a prediction coefficient such that $10<\omega_{n,s}(t^{\prime})<$ 1, $t^{\prime}\leq t$. To achieve accurate predictions, the proposed method leverages the Ordinary Least Squares regression method (OLS) to train these coefficients initially, and use the past experienced observations as the target distribution for training. After that, the current collected data values $V^{\prime}_{n,s}$ and predicted one $\hat{V}_{n,s}$ are shared to all agents via the tree communication structure.
Target. The target $\mathcal{R}$ in the short-term plan selection needs to be iteratively updated to steer the coordination of drones to choose plans for areas and timeslots with abundant sensor data. The percentile-based data filtering is used to eliminate the extreme low sensing values collected by drones. This approach effectively removes the need for data collection in regions with low sensing requirements (e.g., traffic exclusion zones). As a consequence, it helps drones in priortizing their operations in cells and time where sensing values are significantly high. We initialize the target as $r_{n,s}(0)=1$ to encompass all cells and timeslots, and update it as follows:
$\begin{split}r_{n,s}(t)=\left\{\begin{array}[]{ll}0,&V^{\prime}_{n,s}(t)<\overline{v}\wedge p_{n,s}(t)>0\\ r_{n,s}(t-1),&otherwise\end{array},\right.\end{split}$ (10)

### v3

@@ III-B Problem Formulation
$\mathsf{Eff}(a^{u},t)=\frac{\sum^{N}_{n=1}\sum^{S}_{s=1}v_{n,s}(a^{u},t)}{\sum^{N}_{n=1}\sum^{S}_{s=1}V_{n,s}(t)}.$ (3)
@@
$E(a^{u},t)=C^{\mathsf{f}}(u)\cdot t^{\mathsf{f}}(a^{u})+C^{\mathsf{h}}(u)\cdot t^{\mathsf{h}}(a^{u}),$ (5)
where $C^{\mathsf{f}}(u)$ and $C^{\mathsf{h}}(u)$ are the flying and hovering power consumption of drone $u$ respectively; $t^{\mathsf{f}}$ and $t^{\mathsf{h}}$ are the flying and hovering time respectively, which are determined by the sensing plan $P^{u}(a^{u},t)$.
@@
$\mathop{\arg\max}\limits_{a^{u},u\in\mathcal{U}}\;\sum^{U}_{u=1}\sum^{T}_{t=1}[\mathsf{Eff}(a^{u},t)\cdot\mathsf{Acc}(a^{u},t)-E(a^{u},t)].$ (6)
DRL Modeling. Once drones take actions of flying directions, they execute plans and return to charging stations, changing their current locations and battery levels while observing the navigation and sensing of other drones. Therefore, the problem scenario can be explained into a Markov decision process [26]. We model the problem using state, action, and reward concepts: (i) State: The state $\textbf{s}_{t}$ at period $t$ consists of four components $(\mathcal{S}_{1},\mathcal{S}_{2},\mathcal{S}_{3},\mathcal{S}_{4})$, where $\mathcal{S}_{1}=\{c^{u}(t)|_{u\in\mathcal{U}}\}$ is the current locations of drones. Since drones charge at charging stations before taking actions, the location can serve as the index of the charging station. $\mathcal{S}_{2}=\{e^{u}(t)|_{u\in\mathcal{U}}\}$ is the current battery levels of drones, which are calculated based on the battery capacity and energy cost. $\mathcal{S}_{3}=\{P^{u}(t)|_{u\in\mathcal{U}}\}$ is the plan of $u$. $\mathcal{S}_{4}=\{P^{-u}(t)|_{u\in\mathcal{U}}\}$ is the aggregated plan of other drones excluding $u$, which are shared via the designed communication model. (ii) Action: The action $\textbf{a}_{t}=\{a^{1}_{t},...,a^{U}_{t}\}$ at period $t$ consists of period-by-period flying directions (N, E, S, W, NE, SE, SW, NW or O) determined by drones. (iii) Reward: Based on the objective function of Eq.(6), the expected immediate local reward of one drone at period $t$ is defined as follows: $\textbf{r}^{u}_{t}=\mathsf{Eff}(a^{u},t)\cdot\mathsf{Acc}(a^{u},t)-E(a^{u},t)$. Throughout the training process or episodes, the overall reward fluctuates based on the actions taken by the drones. This helps drones in prioritizing areas rich in sensor data by maximizing their reward, i.e., the highest cumulative overall performance.
Figure 1: Overview of the structured tree communication model. During the bottom-up process, Drone 3 aggregates the plans of its children, i.e., drone 1 and drone 2, and sends them to its parent agent with its own plan. During the top-down process, each parent agent sends the total aggregated plans to its children such that all agents obtain the observed plans.
III-C Communication Model
In DRL model, each drone evaluates an immediate reward by considering collective information from other drones. To enhance information exchange efficiency, we design a structured tree communication model. Each drone, controlled by a local agent, connects with other agents into a tree communication structure, as shown in Fig. 1, within which it interacts with its children and parent in a bottom-up and top-down. For example, drones aggregate and share their plans $P^{u}$ in a bottom-up fashion, and then obtain the observed plans $P^{-u}$ in the top-down process [8]. Therefore, this decentralized coordination among agents is highly-efficient with low communication cost [8]. Through iterative bottom-up and top-down exchanges, agents can flexibly change their plans based on the aggregated navigation and sensing from other drones.
Furthermore, agents can be self-organized within the tree to make the system more resilient to attacks or failures [12, 11]. Since agents share the aggregated states instead of the agents’ one, the system is scalable and designed to preserve privacy [^2] Additional privacy protection mechanisms such as differential privacy and homomorphic encryption apply here [34].. For instance, assuming a swarm of drones that belong to various entities, such as aviation agencies, private citizens, and companies, they collaborate to create shared navigation and sensing strategies which rely on the collective sum of plans rather than individual private sensing information.
@@ III-C Communication Model
III-D System Overview
For high-quality sensing performance, the proposed approach DO-RL relies on both decentralized drone coordination for short-term navigation and sensing optimization, as well as long-term scheduling of flying directions to adapt to changing environments. Fig.2 illustrates the designed system framework of DO-RL, consisting of four main components:
Plan generation. This component generates the navigation and sensing options for drones. Given the sensing map, each drone autonomously generates a finite number of discrete plans to initialize the overall process. This provides flexibility for the drones at the next stage to choose in a coordinated way. The generated plans are grouped into different flying directions. This initialization stage is executed only once for each drone.
Plan selection. This distributed component leverages a collective learning to coordinate drones to locally select the optimal navigation and sensing options from their generated plans within a period.
DRL-based scheduling. This component is the core of the framework, which leverages the DRL algorithm to enable drones to take their period-by-period actions of flying directions. Then, the generated subset of plans with the corresponding navigation action are chosen as input for the plan selection. After plan selection, it undergoes two steps: one is periodic state update, which updates the state of drones (i.e., the current location and energy cost) and the system (i.e., the target and required sensing values) for the DRL-based scheduling in the next period; the other is multi-agent reinforcement learning, which is built based on centralized training and decentralized execution.
IV Detailed System Design
Figure 3: Process of path finding in plan generation.
@@ IV-A Plan Generation
Each drone generates nine groups of plans, each associated with an action for a flying direction. These groups consist of multiple individual plans of the same size. A group of plans generated by drone $u$ during a period is defined as $G^{a^{u}}(t)$, where each plan in the group has the same action, $P^{u}(a^{u},t)\in G^{a^{u}}(t)$. Then, a plan can be defined as $P^{u}(a^{u},t)=\{l,K(a^{u}),J(a^{u}),E(a^{u},t)\}$, where $l$ is the index of the plan, $l\leq L$, ($L$ denotes the total number of plans in a group); $K(a^{u})$ denotes the indexes of cells within the searching range along flying direction $a^{u}$; $J(a^{u})$ denotes the indexes of visited cells within $K(a^{u})$; and $E(a^{u},t)$ denotes the energy consumed by $u$ traveling over $J(a^{u})$, i.e., the cost of the plan.
@@
A drone generates a plan by determining its path and calculating the energy required for traversing that path, as shown in Fig. 3. As an example of a plan, the drone $u$ firstly finds a searching range to cells $K(a^{u})$: it determines a range with a certain width (equal to half of distance to nearest charging station) along its flying direction, and then searches the cells within this range. Next, it selects the visited cells $J(a^{u})$ randomly from this range. Then, $u$ finds the shortest possible path among $J(a^{u})$ via the Dijkstra’s algorithm for the Traveling Salesmen Problem. The path includes the information of which cell and which timeslot the drone uses to collect data, and thus both the traveling time $t^{\mathsf{f}}(a^{u})$ and hovering time $t^{\mathsf{h}}(a^{u})$ are calculated. The total energy consumption $E(a^{u},t)$ when $u$ travels over the path is then calculated via Eq.(5).
IV-B Plan Selection
@@
Given the plans in selected flying directions within a period, drones can adapt their assignments by selecting appropriate plans in response to changes in task targets, enhancing their flexibility and adaptability in dynamic environments. In this process, the agents of drones improve their plan selections based on the structured tree communication model. Specifically, each agent obtains the aggregated choices, i.e., the aggregated plan $P^{-u}(a^{u},t)$, from other agents, and chooses one of the plans $P^{u}(a^{u},t,l)$ such that all choices together $P^{u}(a^{u},t)+P^{-u}(a^{u},t)$ add up to match a target $\mathcal{R}(t)$. The target is used to steer each drone to sense over a unique cell during a timeslot, thereby avoiding the simultaneous sensing of multiple drones within the same cell, i.e., preventing over-sensing and under-sensing.
@@
The purpose of coordination is to match the aggregated plans of all agents to the target while minimizing the energy cost $E(a^{u},t)$ of the plan selected by the drone. The cost function for each agent is formulated using the root mean square error (RMSE):
@@
$\begin{split}\mathop{\min}\limits_{a^{u},u\in\mathcal{U}}\;&(1-\beta)\cdot\sqrt{\frac{\sum^{N}_{n=1}\sum^{S}_{s=1}[p_{n,s}(a^{u},t)-r_{n,s}(t)]^{2}}{N\cdot S}}\\ &+\beta\cdot E(a^{u},t),\end{split}$ (7)
where $\beta$ represents the behavior of an agent. As the value of $\beta$ increases, the agent becomes increasingly selfish, prioritizing plans with lower energy costs at the expense of higher root mean squared error. Each agent (or drone) aims to minimize the system-wide cost of sensing while considering its energy consumption. By minimizing this cost function, each drone coordinates to select its optimal plan, indicating a specific path within a period (comprising multiple timeslots). Via planning and selection, drones using DO-RL only need to take high-level actions of flying directions in each period, and thus reducing the number of decisions and observations in DRL.
IV-C Periodic State Update
After choosing a plan, each drone changes its current state at the next period $t+1$, including its location $c^{u}(t+1)$, battery level $e^{u}(t+1)$, selected plan $P^{u}(t+1):=P^{u}(a^{u},t)$, and observed plans $P^{-u}(t+1):=P^{-u}(a^{u},t)$. However, due to dynamic changing environment, drones have no knowledge about the required sensing value in the next period. They need to predict and estimate the required sensing value to calculate reward function via Eq.(3) and (4). Depending on the predicted distribution of sensor data, the target is required to be updated.
Predicted sensing value. The predicted sensing value at period $t$, denoted as $\hat{V}(t)$, is updated in a time-reverse decay, formulated as follows:
$\hat{V}_{n,s}(t)=\sum^{t}_{t^{\prime}=1}(T-t+t^{\prime})\cdot\omega_{n,s}(t^{\prime})\cdot v^{\prime}_{n,s}(a^{u},t^{\prime}):=V_{n,s}(t),$ (8)
where $V^{\prime}_{n,s}$ denotes the data values collected by drones once they execute their selected plans (the total collected values are aggregated and shared to each agent via the top-down interactions in the distributed optimization process [8]); $\omega_{n,s}(t^{\prime})$ is a prediction coefficient such that $10<\omega_{n,s}(t^{\prime})<$ 1, $t^{\prime}\leq t$. To achieve accurate predictions, DO-RL leverages the Ordinary Least Squares regression method (OLS) to train these coefficients initially, and use the past experienced observations as the target distribution for training. After that, the current collected data values $V^{\prime}_{n,s}$ and predicted one $\hat{V}_{n,s}$ are shared to all agents via the tree communication structure.
Target. The target $\mathcal{R}$ in the plan selection needs to be iteratively updated to steer the coordination of drones to choose plans for areas and timeslots with abundant sensor data. The percentile-based data filtering is used to eliminate the extreme low sensing values collected by drones. This approach effectively removes the need for data collection in regions with low sensing requirements (e.g., traffic exclusion zones). As a consequence, it helps drones in priortizing their operations in cells and time where sensing values are significantly high. We initialize the target as $r_{n,s}(0)=1$ to encompass all cells and timeslots, and update it as follows:
$\begin{split}r_{n,s}(t)=\left\{\begin{array}[]{ll}0,&v^{\prime}_{n,s}(a^{u},t)<\overline{v}\wedge p_{n,s}(a^{u},t)>0\\ r_{n,s}(t-1),&otherwise\end{array},\right.\end{split}$ (9)
IV-D Multi-agent Reinforcement Learning
@@
Our approach provides a generalized model for multi-agent reinforcement learning by adopting the framework of centralized training and decentralized execution, as shown in Fig.4. During execution, each agent obtains a local aggregated observation comprising both private and public information. Apart from their own private data such as locations, battery levels, and selected plans, this observation contains aggregated plans shared via the designed communication model (shown in Fig. 1). The training part is built based on an actor-critic model. The actors allow drones to work independently and choose actions based on their observations and policies, allowing for parallel data collection and processing. The critic, however, evaluates the actions of drones and guide them toward collectively optimal decisions. In addition, DO-RL employs Proximal Policy Optimization (PPO) to prevent detrimental updates and improving the stability of the learning process [35].
@@
The process of training primarily involves the following steps: Firstly, each agent takes an action for a flying direction based on the aggregated observation $\textbf{o}^{u}_{t}$ shared via coordination as well as the reward $\textbf{r}^{u}_{t}$. Once all actions are determined, drones select their plans and transition to a new state. Subsequently, the buffer, a data storage structure used for experience replay, stores all transitions of each agent $(\textbf{o}^{u}_{t},\textbf{a}^{u}_{t},\textbf{r}^{u}_{t},\textbf{o}^{u}_{t+1})$. Several groups of transitions $(\textbf{o}^{u}_{k},\textbf{a}^{u}_{k},\textbf{r}^{u}_{k},\textbf{o}^{u}_{k+1})$ are sampled randomly ($H$ groups of transitions) for updating the parameters of both the critic and actor networks. The algorithm is typically an extension of the actor-critic policy gradient approach, employing two deep neural networks for each agent: a critic network $Q(\cdot)$ and an actor network $\pi(\cdot)$. The critic network estimates the reward associated with a transition using Bellman equation, and its parameter $\theta^{Q}$ is then updated by minimizing a loss function $L(\theta^{Q})$:
@@
$\hat{A}^{u}_{k}=\textbf{r}^{u}_{k}+\gamma\cdot Q(\textbf{o}^{u}_{k+1},a^{u}_{k+1})-Q(\textbf{o}^{u}_{k},a^{u}_{k}),$ (11)